OpenStudy (anonymous):

If a number n is subtracted from 25, the result is three less than n. What is the value of n? A. n = 14 B. n = 22 C. n = 28 D. n = 11

OpenStudy (bibby):

Can you translate 3 less than n into an expression?

OpenStudy (bibby):

For reference what we're looking at is a number n is subtracted from 25 = three less than n

OpenStudy (anonymous):

for this problem i got the answer A by using the process of elimination

OpenStudy (anonymous):

I t was A.

OpenStudy (zpupster):

number n is subtracted from 25 25-n three less than n n-3 25-n=n-3 subtract n both sides 25-2n =-3 subtract 25 both sides -2n= -28 divide both sides by -2 n=14 this how it is solved algebracially answer A is confirmed